# RACP: Risk-Aware Contingency Planning with Multi-Modal Predictions

### v1

@@ VI-C4 Branching time estimation
Baseline 4: “Oracle” branching time. This estimator reconstructs the actual branching time by initially simulation the planning problem using a nominal branching time. Subsequently, it extracts the moment of certainty from the retrospective evolution of beliefs. It is important to note that the oracle relies on access to the true human intent and, as such, is not implementable in real-world scenarios. Nevertheless, we incorporate this variant to illustrate the potential performance attainable with the true branching time.
Baseline 5: Branching time heuristics adopted from [63]. This heuristic considers the entropy of the belief that the ego-vehicle maintains over each hypothesis, $\theta\in\Theta$, as an indication of how the observed obstacles’ states are distinct,
@@
Baseline 6: Branching time heuristic adopted from [64]. This heuristic chooses the branching time as the maximum time such that any two future scenarios starting at the current time only diverge by a maximum distance,
$t_{b}=\max_{k\in{2,...,T}}\quad$ $\omega.k$ (21)
@@
TABLE III: Statistical results over 100 experiments for a T-junction scenario with different branching time estimates.
As shown in Tab. III, the performance gap between fixing the branching time to a certain value, and using an oracle estimate is very small. This can be attributed to the fact that since the short-term plan cost is weighted by the belief the ego vehicle maintains over the long plans, the short-term plan tends to be biased toward the long-term plan with the highest belief.
@@ VII Conclusion
where $Q=\texttt{diag}(0,w_{d},w_{v},0)$, and $R=\texttt{diag}(w_{a},w_{\dot{\delta}})$. In this way, we ensure that the cost functions used by the Frenet and branch-MPC planners are similar. Additionally, the same kinematic constraints are applied to all planners including the curvature constraints and the box constraints imposed on the velocity, acceleration, and jerk. Tab. IV summarizes the parameters utilized by the planners in the evaluations.
@@
TABLE IV: List of parameters utilized by the planners for the overtaking scenario evaluations.

### v2

@@ VI-C5 Branching time estimation
Baseline 5: “Oracle” branching time. This estimator reconstructs the actual branching time by initially simulation the planning problem using a nominal branching time. Subsequently, it extracts the moment of certainty from the retrospective evolution of beliefs. It is important to note that the oracle relies on access to the true human intent and, as such, is not implementable in real-world scenarios. Nevertheless, we incorporate this variant to illustrate the potential performance attainable with the true branching time.
Baseline 6: Branching time heuristics adopted from [65]. This heuristic considers the entropy of the belief that the ego-vehicle maintains over each hypothesis, $\theta\in\Theta$, as an indication of how the observed obstacles’ states are distinct,
@@
Baseline 7: Branching time heuristic adopted from [66]. This heuristic chooses the branching time as the maximum time such that any two future scenarios starting at the current time only diverge by a maximum distance,
$t_{b}=\max_{k\in{2,...,T}}$ $\omega.k$ (21)
@@
TABLE IV: Statistical results over 100 experiments for a T-junction scenario with different branching time estimates.
As shown in Tab. IV, the performance gap between fixing the branching time to a certain value, and using an oracle estimate is very small. This can be attributed to the fact that since the short-term plan cost is weighted by the belief the ego vehicle maintains over the long plans, the short-term plan tends to be biased toward the long-term plan with the highest belief.
VI-D Scenario 3. Intersection
In this scenario, the ego-vehicle is tasked with executing a left turn within an urban intersection, all while navigating interactions with multiple obstacle vehicles simultaneously. Each of these obstacle vehicles within the intersection has the option to either yield to the ego-vehicle, thereby allowing it to complete its left turn unimpeded, or to challenge the ego-vehicle and take priority, thereby compelling the ego-vehicle to yield. Similar to the T-junction scenario, we evaluate the efficacy of our proposed approach against established baselines. Quantitative results are presented in Tab V. For consistency, we utilize a horizon of $N=25$ steps, with a discretization step of 0.2 s, resulting in a time horizon of 5.0 s. Additionally, a branching-time of $tb=2.4$ s is defined. Here it should be noted that the prediction model in this scenario is similar to the one employed in the T-junction scenario.
Approach Progress [m] Velocity [m/s] CR
Our method 40.696 (0.582) 8.139 (0.147) 0.00 %
FISS+ [45] 29.225 (0.473) 5.851 (0.168) 0.00 %
Non-cont. MLE [50] 42.261 (0.327) 8.452 (0.263) 3.47 %
Cont. w/o belief updater 37.693 (0.372) 7.531 (0.126) 0.00 %
TABLE V: Statistical results over 100 experiments for the intersection scenario where CR refers to the collision rate.
The outcomes of this scenario mirror those of previous ones, demonstrating that our proposed approach enables the ego-vehicle to successfully execute the left-turn maneuver with enhanced efficiency compared to the robust baseline, where the ego-vehicle is required to yield while waiting for other vehicles to execute their maneuvers inside the intersection [^3] The reader can refer to the video supplement to observe the yielding behavior of the ego-vehicle using the robust baseline.. However, it’s noteworthy that the MLE approach outperforms our method in terms of performance, albeit at the cost of a higher collision rate, as it only considers the most probable mode without accounting for potential deviations.
@@ VII Conclusion
where $Q=\texttt{diag}(0,w_{d},w_{v},0)$, and $R=\texttt{diag}(w_{a},w_{\dot{\delta}})$. In this way, we ensure that the cost functions used by the Frenet and branch-MPC planners are similar. Additionally, the same kinematic constraints are applied to all planners including the curvature constraints and the box constraints imposed on the velocity, acceleration, and jerk. Tab. VI summarizes the parameters utilized by the planners in the evaluations.
@@
TABLE VI: List of parameters utilized by the planners for the overtaking scenario evaluations.
Branching Time Selection
As mentioned in Section VI-C4, the branching time is a critical parameter that affects the contingency planning efficiency. To justify our branching time selection for the simulated scenarios in Section VI, we conducted an ablation study that measures the performance gap between different branching times and an oracle that has access to the true human intent. Fig. 14 illustrates the performance gap regarding the relative velocity for different branching times for both the overtake and T-junction scenarios. As shown, for both scenarios, for the branching times in the middle of the planning horizon, the performance gap compared to the oracle gets smaller. As the branching time gets smaller, a less conservative approach compared to the oracle can be achieved. This comes, however, at the expense of having higher risk as we discussed earlier in Section VI-C4. In contrast, with large branching times, the contingency planning becomes more conservative compared to the oracle. Based on the obtained empirical results, we can conclude that fixing the branching time to a certain value in the middle of the planning horizon can achieve close performance to the updating it based on an oracle in hindsight.
Figure 14: The performance gap regarding the relative velocity for different branching times. Left: overtake scenario, right: T-junction scenario.